OpenStudy (90890908967777878888):

find the number that completes the equation 6/5÷ __ =6

OpenStudy (90890908967777878888):

@misty1212

OpenStudy (taemin):

Any answers to pick from

OpenStudy (misty1212):

HI!!

OpenStudy (90890908967777878888):

hi

OpenStudy (misty1212):

we can do it in our heads

OpenStudy (misty1212):

what would you divide 6 by to get 6?

OpenStudy (90890908967777878888):

1

OpenStudy (misty1212):

right so what would you divide six fifths by to get 6?

OpenStudy (90890908967777878888):

30

OpenStudy (misty1212):

oh no lets go slower

OpenStudy (90890908967777878888):

okay

OpenStudy (misty1212):

\[\frac{6}{5}\div \frac{x}{5}=6\] you want \(x\)

OpenStudy (misty1212):

this is the same as saying \[\frac{6}{5}\times \frac{5}{x}=6\]

OpenStudy (90890908967777878888):

okay

OpenStudy (misty1212):

cancel the fives and get \[\frac{6}{x}=6\]

OpenStudy (90890908967777878888):

1

OpenStudy (taemin):

try this \[6\div6/5 \]

OpenStudy (misty1212):

which is why i asked "what would you divide 6 by to get 6?"

OpenStudy (misty1212):

yes, the answer is 1

OpenStudy (90890908967777878888):

is it 1/5

OpenStudy (misty1212):

six divided by six is one so six fifths divided by one fifth is also 1

OpenStudy (misty1212):

be careful because they are only asking you what goes in the numerator there

OpenStudy (90890908967777878888):

okay

OpenStudy (misty1212):

the numerator is just 1, but \[\frac{6}{5}\div \frac{1}{5}=6\]is correct

OpenStudy (90890908967777878888):

okay

OpenStudy (90890908967777878888):

@misty1212 is it 1/5 or 1/30